IR spectrum 100 TRANSMETTANCES RAVENUMBERScan someone please tell me what groups and/or bonds are present in this IR spectrum?
Answer #1

1100-1120 cm-¹ = saturated alkanes

1700-1725 cm-¹ = carbonyl group

1370-1540 cm-¹ = nitro group

2850-2950 cm-¹ =saturated alkenes